Ticket Number: HEA-NOC/19990302-02 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> UCD
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19990302 09:47 UTC
Ticket Closed: Problem End:
UCD's class B (126.96.36.199) is being held down by the central
exchange routers following flaps on the line.
Traffic between UCD and the NOC was affected for the first ten
minutes of the fault. Traffic between UCD and the internet at large
is still affected.
Continuing to monitor. Ticket will remain open until full routing
davew 19990302 10:37 UTC:
Further flaps at aroudn 10:30 - reseated the BNC connectors at the
NOC to try to reset the line clocking. COntinuing to monitor & will
contact Telecom if further flaps occur. Route remains held down at
Time to Fix:
Not yet known - estimate 10-30 minutes.